# Progress-Aware Video Frame Captioning

### v1

@@ 1 Introduction
Having such progress-aware captions could benefit a great variety of downstream tasks, bringing improved video understanding [78, 73], more precise video retrieval [67, 68, 66], and enriched video generation [58, 50]. Moreover, such a capability could open up new AR/VR and robotics applications. For instance, in AI coaching, a captioning system could meticulously analyze an expert’s tennis forehand, simplifying the learning process for users. Similarly, for how-to video creation, it could elicit and describe the key object state changes at each stage (e.g., “how to make whipped cream”)—useful for both content creators as well as visually-impaired users learning a new skill.
Towards this end, we introduce a novel captioning taskprogress-aware video captioning at the frame level. This task involves generating captions that not only coherently depict action progression but also tailor each description specifically to its corresponding frame. [^2] Without loss of generality, we obtain the input frame sequence by uniformly sampling from the action clips at a fixed rate (1FPS). These frames may or may not demonstrate visual action progression from one to the next, demanding that the model discerns the difference when generating progress-aware captions. Our task is uniquely characterized by its demand for fine-grained temporal sensitivity. By “fine-grained”, we refer to generating detailed, frame-level descriptions that elucidate the stages or procedural steps of the action, effectively conveying how the action is performed throughout the video sequence.
As discussed, traditional video captioning [9, 70, 85, 63] settles for broad event-level descriptions, where a description like scrambling eggs for the video in Figure 1 would be considered entirely accurate. In contrast, we seek progress-aware captions that detail each stage of the action, such as “eggs still intact”, “starting to cook” and “fully mixed”. While recent works [8, 61, 11, 82, 6, 12] enhance the overall descriptiveness of video captions, they continue to produce a single video-level description without distinguishing the nuances between individual frames. Our task delves deeper, exploring how each frame contributes to the narrative of the action’s progression, thereby setting a new standard for fine-grained temporal precision in video captioning.
@@
We propose ProgressCaptioner, a captioning model designed to generate progress-aware frame-level captions. This model is developed through an innovative strategy that interleaves pseudo labeling with two learning stages. In the first stage, we employ multiple VLMs to generate consensus pseudo captions for pairs of frames, and we devise two novel evaluation tasks to automatically assess their quality. In the second stage, armed with that initial model, we apply a two-frame sliding window to generate more precise captions across the full frame sequence, and we generalize our custom evaluation tasks to identify high-quality caption sequences. The final model, ProgressCaptioner, is then trained on this comprehensive dataset, which we term FrameCap, to excel in producing temporally fine-grained captions that accurately characterize action progress.
@@ 2 Related Work
Video Captioning Video captioning [1] aims to produce a single description that encapsulates a video clip. While traditional benchmarks [9, 70, 85, 63] offer a brief one-sentence caption for each video, recent efforts expand this scope, extending captioning to hours-long videos [26], enriching the granularity of details [11, 8, 61], enhancing caption uniqueness [49], integrating a casual temporal narrative [46], or introducing LLM summarization [34].
Adjacent to traditional video captioning, are the tasks of visual storytelling [25, 36] (creating a coherent story for a sequence of snapshots), dense video captioning [30, 74, 86] (temporal localization and captioning of all events in an untrimmed video), and audio description [20, 21, 22] (detailed narrations of visual events in videos (e.g., movies) for visually impaired audiences). However, all these works still address “what is happening” at a coarse-grained event level, e.g., noting that someone is making a souffle within a specific time range. The ability to break down frame-level details—such as whisking egg whites, folding ingredients, and observing the souffle rising—is still lacking.
@@ 3.1 Progress-aware Video Frame Captioning
Problem Formulation Our objective is to develop a captioning model that, given a video, produces accurate and temporally fine-grained captions. Formally, for a sequence of $T$ frames, denoted as $\mathcal{V}=\{v_{i}\}_{i=1}^{T}$, the captioning model generates a corresponding sequence of captions $\mathcal{C}=\{c_{i}\}_{c=1}^{T}$, where each $c_{i}$ describes the $i$-th frame $v_{i}$. This captioning process is underpinned by three key requirements: (1) Accuracy, where each caption $c_{i}$ must faithfully represent what is visually occurring in frame $v_{i}$, without hallucinating from the context of other frames; (2) Temporal Specificity, where each caption $c_{i}$ specifically attends to $v_{i}$, without being overly generic to be applicable to multiple frames in the sequence; (3) Progressive Coherence: The sequence of captions $\{c_{i}\}_{i=1}^{T}$ should build upon each other to reflect the essential changes in the action over time.
@@
Observation I Intuitively, inputting all $T$ frames would seem best. However, current VLM capabilities do not support this extensive context. Specifically, providing too many frames at once often leads to descriptions that lack detail and exhibit temporal inaccuracies, with VLMs also risking memory overload, as similarly observed in [11]. Conversely, providing a single frame at a time reduces the task to image captioning, which is not optimal either, resulting, resul in captions that lack temporal context and coherence.
Figure 3: Captioning outcomes using Gemini-1.5-Pro [53].
Figure 3 shows a representative trial with Gemini-1.5-Pro [53]. Inputting the full sequence (case (a)) yields brief per-frame descriptions with temporal misalignment (i.e., the second caption erroneously describes what is visually occurring in the third frame). On the other hand, captioning frames in isolation (case (b)) removes essential temporal context, where the model mistakes the initial stage of a tennis serve for the follow-through of a forehand swing. These findings underscore the importance of finding a balanced approach and motivate us to adopt a frame pair as the stepping stone of our captioning model development.
@@
Observation II Next, building upon the use of a frame pair ($v_{1},v_{2}$), is the caption pair ($\hat{c}_{1},\hat{c}_{2}$) produced by existing VLMs of sufficient quality to be directly adopted? Our preliminary experiments reveal two main issues: (a) lack of temporal granularity, and (b) temporal hallucination, as showcased in Figure 2. To dissect these issues, we analyze the captions in relation to the visual progression between frames $v_{1}$ and $v_{2}$. Specifically, if there is a visible progression from $v_{1}$ to $v_{2}$ (e.g., the slight peeling of a shrimp’s shell from frame 2 to frame 3 in Figure 2), the captions should adequately reflect this change. Overly similar captions in such a scenario signify a failure in temporal granularity. Conversely, when there is no change between frames (e.g. frames 1 and 2 in Figure 2), the captions should remain consistent. We deem it a temporal hallucination when captions erroneously indicate progression in disagreement with the visuals. This issue often arises in VLMs equipped with strong language decoders, which can introduce a language bias toward generating unwarranted narrative changes.
@@ 3.2 ProgressCaptioner
For pairs passing progression detection, we proceed to our second evaluation task—caption matching—to assess how precisely $\hat{c}_{1}$ and $\hat{c}_{2}$ describe $v_{1}$ and $v_{2}$, respectively. The task is designed as a multi-choice question format, where a VLM is given $\hat{c}_{1}$, $\hat{c}_{2}$, and an “unsure” option, and tasked with matching the correct caption to each frame. A caption pair is considered good if the evaluation VLM correctly identifies $\hat{c}_{1}$ for $v_{1}$ and $\hat{c}_{2}$ for $v_{2}$. Because the captions will all be topically related, this is essentially a matching task with “hard negatives” that lets us automatically gauge the precision of the proposed captions for the target images.
@@
Stage II Training Following the same pipeline as stage I, ProgressCaptioner is first trained through SFT using data prepared during both stages, which includes frame-caption pairs <$\mathbf{v},\mathbf{\hat{c}}^{+}$> and frame-caption sequences <$\mathcal{V},\mathcal{\hat{C}}^{+}$>. Subsequently, we conduct DPO with preference data collected from both stages <$\mathbf{v},\mathbf{\hat{c}}^{+},\mathbf{\hat{c}}^{-}$> and <$\mathcal{V},\mathcal{\hat{C}}^{+},\mathcal{\hat{C}}^{-}$>  to further refine model performance and mitigate hallucination. This sequential approach results in our final captioning model. The framework is illustrated in Figure 4.
@@ 4 Experiments
Table 1: Results on the FrameCapEval Benchmark, composed of video from four public datasets. Cap and Prog denote caption matching and progression detection accuracy, respectively. ProgressCaptioner greatly outperforms SOTA open-source VLMs and even the leading proprietary models, despite being a 7B model. The best results are bolded and underlined, the second best are bolded, and the third best are italicized. Moreover, the results confirm our model’s generalizability from in-domain datasets (HTC for HowToChange and COIN) to external datasets not seen during training (Penn&K, representing Penn Action and Kinetics).
@@ 4.1 Benchmarking Video Frame Captioning
To systematically evaluate the quality of generated video frame-wise captions, we propose the FrameCapEval benchmark. Our benchmark sets a new standard for temporal precision in video captioning, assessing captions based on the criteria established in Sec. 3.1.
Evaluation Metrics We employ the automatic evaluation tasks of progression detection and caption matching (Sec. 3.2), reporting accuracy with Llama-3.1-70B-Instruct [14] as the evaluation LLM and Gemini-1.5-Pro [53] as the evaluation VLM (see Supp. for details). Additionally, we enhance our evaluation with a user study of 15 participants, reporting the average selection rate.
@@
Baselines We evaluate an array of state-of-the-art VLMs, including two proprietary models, GPT-4o [2] and Gemini-1.5-Pro [53], and five open-source models—Idefics2 [31], VILA [40], Qwen2-VL [62], LLAVA-Video [82], and LLAVA-OV [33], all of which process the entire $T$-frame sequences as input. To ensure a comprehensive evaluation, we also incorporate image captioning baselines using Gemini-1.5-Pro and LLAVA-OV. We select open-source VLM variants with fewer than 10B parameters for computational efficiency and a fair comparison with our ProgressCaptioner, which is an 7B model. The closed proprietary models are much larger and trained with much more extensive data; we include them as a useful reference point, but stress that they do not constitute an apples-to-apples comparison, to the disadvantage of our ProgressCaptioner.
Implementation ProgressCaptioner is constructed with SigLIP [79] as the vision encoder and Qwen2 [62] as the language model, linked through a projector, and initialized from the LLAVA-OV-7B checkpoint [33]. For final-stage model training, we finetune ProgressCaptioner on 274.5K data samples over 1 epoch during the SFT stage and on 47.3K preference data for 1 epoch during the DPO stage.
@@ 4.2 Applications of Video Frame Captioning
Keyframe Selection Our first use-case leverages ProgressCaptioner’s temporally precise captions as an intermediate representation to identify keyframes within a densely sampled sequence, aided by an LLM (see Supp for details). Figure 7 provides two examples, showcasing how ProgressCaptioner’s produced captions allow selecting distinct frames that effectively capture different stages of the welding and squeezing grapefruit action. While recent video summarization work [24] explores using VLMs and LLMs for keyframe selection, it aims to identify coarse-grained events within long videos, which is not adequate for our problem scenario. See Supp. for a side-to-side comparison and more qualitatives that underscore this distinction.
@@
Zero-shot Frame Classification We repurpose zero-shot frame-wise classification task into a multi-choice QA format, using frame-wise captions to guide an LLM in identifying the correct label per frame, evaluating caption accuracy and granularity (Figure 9 left). Results (Figure 9 right) show that ProgressCaptioner consistently outperforms LLAVA-OV across both datasets. Notably, our training involves no signals related to these frame-wise labels, underscoring its generalizability and effectiveness in enhancing video frame-level understanding.
@@
Video QA Finally, we report results using frame-wise descriptions for video QA, where an LLM (we use GPT-4o) is employed to answer questions on NExT-QA (ATP-Hard) set. As shown in Table 2, ProgressCaptioner achieves the best results on this benchmark, outperforming the previous leader VideoAgent [64] by +3.4%. Compared with a similar setup using LLAVA-OV, ProgressCaptioner achieves a +4.7% gain in the temporal subset, highlighting its superior ability to produce fine-grained, temporally precise descriptions and bring enhanced video understanding.
@@ 5 Conclusion
We introduce progress-aware video frame captioning, which necessitates a significant enhancement in current captioning models’ capability to describe temporal action progression precisely. Towards this end, we develop ProgressCaptioner and show its effectiveness in enhancing the temporal precision and alignment of captions with corresponding frames. Furthermore, we demonstrate its practical applications: keyframe selection and enhanced video understanding. By setting a new standard for temporal precision in video captioning, we hope our work inspires further development in this evolving domain.
@@ 1.1 FrameCap Training Data
For each frame sequence, the caption sequence that passes and fails these evaluations forms our preference data, which is utilized for DPO training of ProgressCaptioner. See Figure 10 and  11 for examples of frame pair data obtained from progression detection and caption matching, respectively, and Figure 12 for an illustration of the frame sequence data preparation process. Table 3 provides a summary of the training data statistics. The first data preparation stage collects a total of 240K frame-caption pairs for supervised fine-tuning (SFT) and 21K preference pairs for direct preference optimization (DPO). The second stage further expands the dataset to include 34K multi-frame and caption sequences for SFT, along with 26K frame-caption sequences for DPO.

### v2

@@ 1 Introduction
Having such progress-aware captions could benefit a great variety of downstream tasks, bringing improved video understanding [79, 73], more precise video retrieval [67, 68, 66], and enriched video generation [58, 50]. Moreover, such a capability could open up new AR/VR and robotics applications. For instance, in AI coaching, a captioning system could meticulously analyze an expert’s tennis forehand, simplifying the learning process for users. Similarly, for how-to video creation, it could elicit and describe the key object state changes at each stage (e.g., “how to make whipped cream")—useful for both content creators as well as visually-impaired users learning a new skill. See Fig. 2.
Towards this end, we introduce a novel captioning taskprogress-aware video captioning at the frame level. This task involves generating captions that not only coherently depict action progression but also tailor each description specifically to its corresponding frame. [^3] Without loss of generality, we obtain the input frame sequence by uniformly sampling from the action clips at a fixed rate (1FPS). These frames may or may not demonstrate visual action progression from one to the next, demanding that the model discern the difference when generating progress-aware captions. Our task is uniquely characterized by its demand for fine-grained temporal sensitivity. By “fine-grained”, we refer to generating detailed descriptions that elucidate the stages or procedural steps of the action, effectively conveying how the action is performed throughout the video sequence.
As discussed, traditional video captioning [9, 70, 86, 63] settles for broad event-level descriptions, where a description like scrambling eggs for the video in Figure 1 would be considered entirely accurate. In contrast, we seek progress-aware captions that detail each stage of the action, such as “eggs still intact”, “starting to cook” and “fully mixed”. While recent works [8, 61, 11, 83, 6, 12] enhance the overall descriptiveness of video captions, they continue to produce a single video-level description without distinguishing the nuances across time. Our task delves deeper, exploring how each frame contributes to the narrative of the action’s progression, thereby setting a new standard for fine-grained temporal precision in video captioning.
Figure 2: Use cases of video frame captioning: finer-grained captions enable detailed, step-by-step guidance for daily tasks.
@@
We introduce ProgressCaptioner, a model for generating progress-aware frame-level video captions. Our approach uniquely interleaves pseudo labeling with two learning stages. Stage I develops a frame pair captioning model, and stage II extends this to full frame sequences. This process also creates the FrameCap training dataset, comprising action videos along with high-quality frame-wise captions, which are initially generated by an ensemble of VLMs and then filtered using our proposed evaluation methods.
@@ 2 Related Work
Video Captioning Video captioning [1] aims to produce a single description that encapsulates a video clip. While traditional benchmarks [9, 70, 86, 63] offer a brief one-sentence caption for each video, recent efforts expand this scope, extending captioning to hours-long videos [26], enriching the granularity of details [11, 8, 61], enhancing caption uniqueness [49], integrating a causal temporal narrative [46], or introducing LLM summarization [34].
Adjacent to traditional video captioning are the tasks of visual storytelling [25, 36] (creating a coherent story for a sequence of snapshots), dense video captioning [30, 74, 87] (temporal localization and captioning of all events in an untrimmed video), audio description [20, 21, 22] (detailed narrations of visual events in videos (e.g., movies) for visually impaired audiences), and video paragraph captioning [78] (producing a multi-sentence paragraph describing the video). However, all these works still address “what is happening” at a coarse-grained event level, e.g., noting that someone is making a souffle within a specific time range. The ability to break down frame-level details—such as whisking egg whites, folding ingredients, and observing the souffle rising—is still lacking.
@@ 3.1 Progress-aware Video Frame Captioning
Problem Formulation Our objective is to develop a captioning model that, given a video, produces accurate temporally fine-grained captions. Formally, for a sequence of $T$ frames, denoted as $\mathcal{V}=\{v_{i}\}_{i=1}^{T}$, the captioning model generates a corresponding sequence of captions $\mathcal{C}=\{c_{i}\}_{c=1}^{T}$, where each $c_{i}$ describes the $i$-th frame $v_{i}$ (recall we sample at 1FPS). This captioning process has three key requirements: (1) Accuracy, where each caption $c_{i}$ must faithfully represent what is visually occurring in frame $v_{i}$, without hallucinating from the context of other frames; (2) Temporal Specificity, where each caption $c_{i}$ specifically attends to $v_{i}$, without being overly generic to be applicable to multiple frames in the sequence; (3) Progressive Coherence: The sequence of captions $\{c_{i}\}_{i=1}^{T}$ should build upon each other to reflect the essential changes in the action over time.
@@
Observation I Intuitively, inputting all $T$ frames would seem best. However, current VLM capabilities do not support this extensive context. Specifically, providing too many frames at once often leads to descriptions that lack detail and exhibit temporal inaccuracies, with VLMs also risking memory overload, as similarly observed in [11]. Conversely, providing a single frame at a time reduces the task to image captioning, which is not optimal either, resulting in captions that lack temporal context and coherence.
Figure 4 shows a representative trial with Gemini-1.5-Pro [53]. Inputting the full sequence (case (a)) yields brief per-frame descriptions with temporal misalignment (i.e., the second caption erroneously describes what is visually occurring in the third frame). On the other hand, captioning frames in isolation (case (b)) removes essential temporal context, where the model mistakes the initial stage of a tennis serve for the follow-through of a forehand swing. These findings underscore the importance of finding a balanced approach and motivate us to adopt a frame pair as the stepping stone of our captioning model.
@@
Observation II Next, building upon the use of a frame pair ($v_{1},v_{2}$), is the caption pair ($\hat{c}_{1},\hat{c}_{2}$) produced by existing VLMs of sufficient quality to be directly adopted? Our preliminary experiments reveal two main issues: (a) lack of temporal granularity, and (b) temporal hallucination, as showcased in Figure 3. To dissect these issues, we analyze the captions in relation to the visual progression between frames $v_{1}$ and $v_{2}$. Specifically, if there is a visible progression from $v_{1}$ to $v_{2}$ (e.g., the slight peeling of a shrimp’s shell from frame 2 to frame 3 in Figure 3), the captions should adequately reflect this change. Overly similar captions in such a scenario signify a failure in temporal granularity. Conversely, when there is no change between frames (e.g. frames 1 and 2 in Figure 3), the captions should remain consistent. We deem it a temporal hallucination when captions erroneously indicate progression in disagreement with the visuals. This deficiency in existing VLMs motivates our development of a new captioning model and specialized evaluations for high-quality caption selection.
@@ 3.2 ProgressCaptioner
For pairs passing progression detection, we proceed to our second evaluation task—caption matching—to assess how precisely $\hat{c}_{1}$ and $\hat{c}_{2}$ describe $v_{1}$ and $v_{2}$, respectively. The task is designed as a multi-choice question format, where a VLM is given $\hat{c}_{1}$, $\hat{c}_{2}$, and an “unsure” option, and tasked with matching the correct caption to each frame. A caption pair is considered good if the evaluation VLM correctly identifies $\hat{c}_{1}$ for $v_{1}$ and $\hat{c}_{2}$ for $v_{2}$. Because the captions will all be topically related, this is essentially a matching task with “hard negatives" that lets us automatically gauge the precision of the proposed captions for the target images.
@@
Stage II Training Following the same pipeline as stage I, ProgressCaptioner is first trained through SFT using data prepared during both stages, which includes frame-caption pairs <$\mathbf{v},\mathbf{\hat{c}}^{+}$> and frame-caption sequences <$\mathcal{V},\mathcal{\hat{C}}^{+}$>. Subsequently, we conduct DPO with preference data collected from both stages <$\mathbf{v},\mathbf{\hat{c}}^{+},\mathbf{\hat{c}}^{-}$> and <$\mathcal{V},\mathcal{\hat{C}}^{+},\mathcal{\hat{C}}^{-}$>  to further refine performance and mitigate hallucination. This sequential approach results in our final captioning model, that accepts inputs ranging from 2 to $T$ frames. This flexibility allows users to control the temporal context, balancing the need for local frame-wise changes (smaller window) and global event progressions (larger window). The framework is illustrated in Figure 5.
@@ 4 Experiments
Table 1: Results on the FrameCapEval Benchmark, composed of video from four public datasets. Cap and Prog denote caption matching and progression detection accuracy, respectively. PL denotes the pseudo labeling baseline adopting filtered captions from multiple VLMs. ProgressCaptioner greatly outperforms SOTA open-source VLMs and even the leading proprietary models, despite being a 7B model. The best results are bolded and underlined, the second best are bolded, and the third best are italicized. The results confirm our model’s generalizability from in-domain datasets (HTC for HowToChange and COIN) to external datasets not seen during training (Penn&K for Penn Action and Kinetics).
@@ 4.1 Benchmarking Video Frame Captioning
Evaluation Metrics We employ the automatic evaluation tasks of progression detection and caption matching (Sec. 3.2), reporting accuracy with Llama-3.1-70B-Instruct [14] as the evaluation LLM and Gemini-1.5-Pro [53] as the evaluation VLM. Additionally, we enhance our evaluation with a user study of 15 participants, reporting the average selection rate. See Supp. for experiments on evaluation metric reliability and full user study details.
Baselines We evaluate an array of state-of-the-art VLMs, including two proprietary models, GPT-4o [2] and Gemini-1.5-Pro [53], and five open-source models—Idefics2 [31], VILA [40], Qwen2-VL [62], LLAVA-Video [83], and LLAVA-OV [33]. We also include a pseudo labeling baseline using filtered captions produced by an ensemble of VLMs (Sec. 3), and image captioning baselines using Gemini-1.5-Pro and LLAVA-OV. We select open-source VLM variants with fewer than 10B parameters for computational efficiency and a fair comparison with our ProgressCaptioner, which is an 7B model. The closed proprietary models are much larger and trained with much more extensive data; we include them as a useful reference point, but stress that they do not constitute an apples-to-apples comparison, to the disadvantage of our ProgressCaptioner.
Implementation ProgressCaptioner is constructed with SigLIP [80] as the vision encoder and Qwen2 [62] as the language model, linked through a projector, and initialized from the LLAVA-OV-7B checkpoint [33]. For benchmark evaluations in Sec. 4.1, ProgressCaptioner operates on the full $T$-frame sequence for a comprehensive temporal context. For applications presented in Sec. 4.2, where fine-grained analysis of local frame changes is crucial, a sliding window approach is used, with the model processing frame pairs. To ensure a fair comparison, all video baseline models are provided with the same temporal context as ProgressCaptioner across all evaluations.
@@ 4.2 Applications of Video Frame Captioning
Keyframe Selection Our first use-case leverages ProgressCaptioner’s temporally precise captions as an intermediate representation to identify keyframes within a densely sampled sequence, aided by an LLM (see Supp. for details). Figure 8 provides two examples, showcasing how ProgressCaptioner’s produced captions allow selecting distinct frames that effectively capture different stages of the welding and squeezing grapefruit action. While recent video summarization work [24] explores using VLMs and LLMs for keyframe selection, it aims to identify coarse-grained events within long videos, which is not adequate for our problem scenario. See Supp. for a side-to-side comparison and more qualitatives that underscore this distinction.
@@
(a) Zero-shot Frame Classification We repurpose zero-shot frame-wise classification task into a multi-choice QA format, using frame-wise captions to guide an LLM in identifying the correct label per frame, evaluating caption accuracy and granularity (Figure 10 left). Results (Figure 10 right) show that ProgressCaptioner consistently outperforms LLAVA-OV across both datasets. Notably, our training involves no signals related to these frame-wise labels, underscoring its generalizability and effectiveness in enhancing video frame-level understanding.
@@
(b) Video QA Finally, we report results using frame-wise descriptions for video QA, where an LLM (we use GPT-4o) is employed to answer questions on NExT-QA (ATP-Hard) set. As shown in Table 2, ProgressCaptioner achieves the best results on this benchmark, outperforming the previous leader VideoAgent [64] by +3.4%. Compared with a similar setup using LLAVA-OV, ProgressCaptioner achieves a +4.7% gain in the temporal subset, highlighting its superior ability to produce fine-grained, temporally precise descriptions and bring enhanced video understanding.
@@ 5 Conclusion
We introduce progress-aware video frame captioning, which necessitates a significant enhancement in current captioning models’ capability to describe temporal action dynamics. Towards this end, we develop ProgressCaptioner and show its effectiveness in enhancing the temporal precision and alignment of captions with corresponding frames. Furthermore, we demonstrate its practical applications: keyframe selection and enhanced video understanding. By setting a new standard for temporal precision in video captioning, we hope our work inspires further development in this domain.
Acknowledgements
Research supported in part by the IFML NSF AI Institute.
@@ 1.1 FrameCap Training Data
For each frame sequence, the caption sequence that passes and fails these evaluations forms our preference data, which is utilized for DPO training of ProgressCaptioner. See Figure 11 and  12 for examples of frame pair data obtained from progression detection and caption matching, respectively, and Figure 13 for an illustration of the frame sequence data preparation process. Table 3 provides a summary of the training data statistics. The first data preparation stage collects a total of 240K frame-caption pairs for supervised fine-tuning (SFT) and 21K preference pairs for direct preference optimization (DPO). The second stage further expands the dataset to include 34K multi-frame and caption sequences for SFT, along with 26K frame-caption sequences for DPO.